# Meta-Semantics Augmented Few-Shot Relational Learning

### v3

@@ Abstract
Han Wu1,2and Jie Yin1 1The University of Sydney, Australia 2Peking University, China {han.wu, jie.yin@sydney.edu.au}
@@ 1 Introduction
To fill this gap, we propose PromptMeta, a novel prompted meta-learning framework for few-shot relational learning. PromptMeta integrates meta-semantic knowledge with relational information inherent in KGs to improve generalization to rare or new relations. Inspired by the recent success of prompting techniques in NLP Zhou et al. (2022); ENG , PromptMeta incorporates prompt learning into meta-representation learning for few-shot tasks, offering a distinctive approach to improving generalization. Its novelty lies in two aspects: First, we propose a Meta-Semantic Prompt (MSP) pool that dynamically learns and consolidates high-level meta-semantics shared across tasks, facilitating effective knowledge transfer. Second, we introduce a learnable fusion mechanism that effectively integrates these meta-semantics with task-specific relational information, tailored to individual tasks. Both components are jointly optimized with model parameters within a meta-learning framework. Our main contributions are as follows:
@@ 2 Related Works
Built upon the recent success in language model fine-tuning, prompting techniques are increasingly explored for tasks that combine large language models (LLMs) with structured knowledge sources such as graphs Yu et al. (2022). Hard text prompts, commonly used in NLP tasks Zhou et al. (2022); ENG , prepend handcrafted instructions to an input text, guiding models to extract relevant task-relevant knowledge using masked language modeling for downstream tasks. For example, G-Prompt (Huang et al., 2023) uses hard text prompts to extract task-relevant node features, which are then fed into a GNN layer for few-shot node classification. TAPE He et al. (2024) prompts an LLM to retrieve textual explanations as features to enrich node representations for downstream GNN training. Alternatively, learnable prompts have been proposed to reduce the high engineering costs in prompt design. A common approach is to define a graph template shared by both pre-training and downstream tasks for better adaptation, often coupled with specific GNN layers. GraphPrompt Liu et al. (2023) and HGPrompt Yu et al. (2024) are exemplars that integrate prompts as learnable vectors within a hidden GNN layer by modifying the Readout function or its input.
@@ 4.2 Learnable Meta-Semantic Prompt Pool
$\mathcal{L}_{\mathrm{pt}}=\frac{1}{K}\sum_{i=1}^{K}-\log\frac{\exp\left(\eta(\mathbf{r}^{(p)},[\mathbf{h}^{(s)}_{i};\mathbf{t}^{(s)}_{i}])/\tau\right)}{\sum^{N}_{n=1}\exp\left(\eta(\mathbf{r}^{(p)},{\mathbf{r}^{-}}^{(p)}_{n})/\tau\right)},$ (6)
@@ 4.4 Model Optimization
[…]
where $||\cdot||_{2}$ denotes the $\ell_{2}$ norm.
To effectively learn meta-representations, we minimize a margin-based ranking loss over the support set:
$\mathcal{L}(\mathcal{S}_{r})=\sum_{(h_{i},t_{i})\in\mathcal{S}_{r}}\max\{0,\mathrm{score}(h_{i},t_{i})+\gamma-\mathrm{score}(h_{i},t_{i}^{-})\},$ (10)
@@
[…]
@@
$\mathcal{L}(\mathcal{Q}_{r})=\sum_{(h_{j},t_{j})\in\mathcal{Q}_{r}}\max\{0,\mathrm{score}(h_{j},t_{j})+\gamma-\mathrm{score}(h_{j},t_{j}^{-})\},$ (13)
$(h_{j},r,t_{j}^{-})$ is a negative triplet generated similar to $(h_{i},r,t_{i}^{-})\in\mathcal{S}_{r}$.
@@ 5.7 Case Study
Lastly, we present a case study to visually illustrate the efficacy of PromptMeta in augmenting relational embeddings with meta-semantic information for FSRL. Figure 2 shows the t-SNE visualization of three relations from Nell-One: color_of_object, person_born_in_location, and person_die_in_country, under the 5-shot setting. In the figure, dotted arrows indicate relational embeddings conventionally used in FSRL, whereas solid arrows denote meta-semantic prompt (MSP) vectors learned by PromptMeta. As expected, the MSP vectors for the semantically related relations person_born_in_location and person_die_in_country exhibit greater similarity, as evidenced by a smaller angle $\theta_{2}$ between them. This angle is notably smaller than $\theta_{3}$, the angle formed by the MSP vectors of the dissimilar pair person_die_in_country and color_of_object. This observation affirms the ability of PromptMeta to learn more generalizable semantic patterns shared across similar relations. By incorporating the learned MSP vectors to augment relational embeddings, PromptMeta effectively reduces the angular distance ($\theta_{1}$) between the embeddings of semantically related relations, thereby improving generalization across few-shot relations. Refer to Appendix C for a quantitative case study on the benefits of utilizing semantic information for FSRL.

### v4

@@ Abstract
Han Wu1,2and Jie Yin1thanks: Corresponding author 1The University of Sydney, Australia 2Peking University, China {han.wu, jie.yin@sydney.edu.au}
@@ 1 Introduction
To fill this gap, we propose PromptMeta, a novel prompted meta-learning framework for few-shot relational learning. PromptMeta integrates meta-semantic knowledge with relational information inherent in KGs to improve generalization to rare or new relations. Inspired by the recent success of prompting techniques in NLP Zhou et al. (2022); Yu et al. (2025), PromptMeta incorporates prompt learning into meta-representation learning for few-shot tasks, offering a distinctive approach to improving generalization. Its novelty lies in two aspects: First, we propose a Meta-Semantic Prompt (MSP) pool that dynamically learns and consolidates high-level meta-semantics shared across tasks, facilitating effective knowledge transfer. Second, we introduce a learnable fusion mechanism that effectively integrates these meta-semantics with task-specific relational information, tailored to individual tasks. Both components are jointly optimized with model parameters within a meta-learning framework. Our main contributions are as follows:
@@ 2 Related Works
Built upon the recent success in language model fine-tuning, prompting techniques are increasingly explored for tasks that combine large language models (LLMs) with structured knowledge sources such as graphs Yu et al. (2022). Hard text prompts, commonly used in NLP tasks Zhou et al. (2022); Yu et al. (2025), prepend handcrafted instructions to an input text, guiding models to extract relevant task-relevant knowledge using masked language modeling for downstream tasks. For example, G-Prompt (Huang et al., 2023) uses hard text prompts to extract task-relevant node features, which are then fed into a GNN layer for few-shot node classification. TAPE He et al. (2024) prompts an LLM to retrieve textual explanations as features to enrich node representations for downstream GNN training. Alternatively, learnable prompts have been proposed to reduce the high engineering costs in prompt design. A common approach is to define a graph template shared by both pre-training and downstream tasks for better adaptation, often coupled with specific GNN layers. GraphPrompt Liu et al. (2023) and HGPrompt Yu et al. (2024) are exemplars that integrate prompts as learnable vectors within a hidden GNN layer by modifying the Readout function or its input.
@@ 4.2 Learnable Meta-Semantic Prompt Pool
$\mathcal{L}_{\mathrm{pt}}=\frac{1}{K}\sum_{i=1}^{K}-\log\frac{\exp\left(\eta(\mathbf{r}^{(p)},[\mathbf{h}^{(s)}_{i};\mathbf{t}^{(s)}_{i}])/\tau\right)}{\sum\limits^{N}_{n=1}\exp\left(\eta(\mathbf{r}^{(p)},{\mathbf{r}^{-}}^{(p)}_{n})/\tau\right)},$ (6)
@@ 4.4 Model Optimization
$\mathrm{score}(h_{i},t_{i})=\|\mathbf{h}^{\prime}_{i}+\mathbf{r}^{m}-\mathbf{t}^{\prime}_{i}\|_{2},$ (9)
where $\|\cdot\|_{2}$ denotes the $\ell_{2}$ norm.
To learn meta-representations, we minimize a margin-based ranking loss over the support set:
$\mathcal{L}(\mathcal{S}_{r})=\sum_{\mathclap{(h_{i},r,t_{i})\in\mathcal{S}_{r}}}\max\big\{0,\mathrm{score}(h_{i},t_{i})+\gamma-\mathrm{score}(h_{i},t_{i}^{-})\big\},$ (10)
@@
$\mathrm{score}(h_{j},t_{j})$ $=\|\textbf{h}^{\prime}_{j}+\mathbf{r}^{m^{\prime}}-\textbf{t}^{\prime}_{j}\|_{2}.$ (12)
@@
$\mathcal{L}(\mathcal{Q}_{r})=\sum_{\mathclap{(h_{j},r,t_{j})\in\mathcal{Q}_{r}}}\max\big\{0,\mathrm{score}(h_{j},t_{j})+\gamma-\mathrm{score}(h_{j},t_{j}^{-})\big\},$ (13)
$(h_{j},r,t_{j}^{-})$ is a negative triplet generated similarly to $(h_{i},r,t_{i}^{-})\in\mathcal{S}_{r}$.
@@ 5.7 Case Study
Lastly, we present a case study to visually illustrate the efficacy of PromptMeta in augmenting relational embeddings with meta-semantic information for FSRL. Figure 2 shows the t-SNE visualization of three relations from Nell-One: color_of_object, person_born_in_location, and person_die_in_country, under the 5-shot setting. In the figure, dotted arrows indicate relational embeddings conventionally used in FSRL, whereas solid arrows denote meta-semantic prompt (MSP) vectors learned by PromptMeta. As expected, the MSP vectors for the semantically related relations person_born_in_location and person_die_in_country exhibit greater similarity, as evidenced by a smaller angle $\theta_{2}$ between them. This angle is notably smaller than $\theta_{3}$, the angle formed by the MSP vectors of the dissimilar pair person_die_in_country and color_of_object. This observation affirms the ability of PromptMeta to learn more generalizable semantic patterns shared across similar relations. By incorporating the learned MSP vectors to augment relational embeddings, PromptMeta effectively reduces the angular distance ($\theta_{1}$) between the embeddings of semantically related relations, thereby improving generalization across few-shot relations. Please refer to Appendix C for a quantitative case study on the benefits of utilizing semantic information for FSRL.